# CleAR: Robust Context-Guided Generative Lighting Estimation for Mobile Augmented Reality

### v1

@@ Abstract.
High-quality environment lighting is the foundation of creating immersive user experiences in mobile augmented reality (AR) applications. However, achieving visually coherent environment lighting estimation for Mobile AR is challenging due to several key limitations associated with AR device sensing capabilities, including limitations in device camera FoV and pixel dynamic ranges. Recent advancements in generative AI, which can generate high-quality images from different types of prompts, including texts and images, present a potential solution for high-quality lighting estimation. Still, to effectively use generative image diffusion models, we must address their key limitations of generation hallucination and slow inference process. To do so, in this work, we design and implement a generative lighting estimation system called CleAR that can produce high-quality and diverse environment maps in the format of $360^{\circ}$ images. Specifically, we design a two-step generation pipeline guided by AR environment context data to ensure the results follow physical environment visual context and color appearances. To improve the estimation robustness under different lighting conditions, we design a real-time refinement component to adjust lighting estimation results on AR devices. To train and test our generative models, we curate a large-scale environment lighting estimation dataset with diverse lighting conditions. Through quantitative evaluation and user study, we show that CleAR outperforms state-of-the-art lighting estimation methods on both estimation accuracy and robustness. In particular, CleAR achieves 56% - 51% (an average of 53%) accuracy improvement on virtual object renderings across objects with three distinctive types of materials over different reflective properties. Moreover, CleAR supports real-time refinement of lighting estimation results, ensuring robust and timely environment lighting updates for AR applications. Our end-to-end generative estimation takes as fast as 3.2 seconds, outperforming state-of-the-art methods by 110$\times$.
@@ 1. Introduction
We define the lighting estimation task for AR systems as estimating a complete environment map (a 360 HDR image) from partial observation of the environment (an LDR image with limited FoV) in real-time. This task is critical for supporting three key aspects of photorealistic rendering in real-time AR systems. (i) Reflective object rendering, which requires a complete environment map with visually coherent details. (ii) Highlights and shadows rendering, which requires HDR pixel information in environment maps. (iii) Temporally consistent visual coherency, which requires the system to robustly adapt to changing environmental lighting in real-time.
To address this challenging task, traditional systems adopt autoregressive models (Zhao and Guo, 2021). These models extract low-frequency information based on camera input (shown in Fig. 1a). While these methods are able to capture coarse information about environmental lighting conditions, high-frequency details are missing. These details are important for creating a visually coherent AR system. To obtain the details, recent works (Somanath and Kurz, 2021; Yang et al., 2023; Wang et al., 2022) leverage the advancement of controllable generative models. These models have the potential to extract fine-grained environment details (shown in Fig. 1a), empowering AR systems with a better rendering effect. While generative-model-based methods seem promising, we identify two key challenges in using them on mobile AR systems.
@@
First, robust lighting estimation demands accurate estimation under challenging lighting conditions. However, through a measurement study on existing datasets, we found that existing lighting estimation datasets have several distribution biases in key lighting properties, including light intensity and color temperatures. In further tests, we found that these data biases will affect the generalization and robustness of multiple recent lighting estimation models (Wang et al., 2022; Akimoto et al., 2022). Therefore, data diversity and fairness must be ensured when training and testing generative lighting estimation systems. Second, interactive mobile AR applications demand timely updates of lighting estimation results. However, generative models usually experience long inference latency. Without system latency optimization, naively integrating generative lighting estimation models into AR applications can easily violate the temporal constituency of immersive user experiences.
@@
In this paper, we address the above issues with CleAR, an edge-assisted novel generative lighting estimation system for mobile AR. To tackle AR device limitations in observation capability, we first design a two-step generative pipeline that estimates 360 HDR environment maps from partial LDR environment observations. Our key design insight is to separate the generative model training objective into two domains: LDR environment map completion and high-intensity pixel value estimation. This novel learning objective design addresses the practical challenges of the scarcity of high-quality lighting estimation training data, and allows us to leverage pre-trained large model to tackle each generation step effectively.
On top of the two-stepped generative pipeline, CleAR uses controlling signals extracted from AR context data to ensure the estimated environment map aligns with the environment of physical AR devices. Specifically, CleAR uses environment semantic maps, which control the environment map visual details, and device ambient light sensor data, which informs the lighting intensity and color temperatures during generation. Specifically, this information is used as image and text inputs for ControlNet models. This design helps the generative pipeline to handle the significant information increases during environment map generation.
Generative model inference latency is high, even on the edge servers. Therefore, we design an edge-device collaborative estimation system architecture with on-device refinement components to adjust edge estimation results to real-time lighting conditions. Specifically, we design a multi-output estimation strategy with an on-device adaptive output selection component. The estimation strategy is configured with the optimal number of generation outputs based on our performance measurement results. Additionally, CleAR uses a color appearance matching technique to efficiently and effectively adjust edge estimation results, even for challenging environment lighting conditions.
@@
To evaluate the lighting estimation quality of CleAR, we integrate CleAR into a mobile AR application using Unity, Python, and a recent AR data streaming framework ARFlow (Zhao and Guo, 2024). We compare CleAR’s virtual object rendering quality to ones using environment lighting acquired from three representative baselines: unwrapping a mirror ball (physical reference) (Debevec, 2006), ARKit (Apple, 2022) (commercial), and LitAR (Zhao et al., 2022) (academic). Our evaluation shows that CleAR can generate environment maps with better image details to support more visually coherent virtual object rendering. To quantitatively understand CleAR’s performance, we evaluate CleAR with standard testing dataset (Phongthawee et al., 2023) and compare with state-of-the-art lighting estimation models (Phongthawee et al., 2023; Wang et al., 2022; Akimoto et al., 2022; Gardner et al., 2017). Our evaluation shows that CleAR outperforms DiffusionLight (Phongthawee et al., 2023) by up to 56% using the three-sphere evaluation protocol (Wang et al., 2022). Also, notably, CleAR achieves 110X estimation latency reduction compared to DiffusionLight due to our efficient generative pipeline and generation control design. Furthermore, we verify the robustness of CleAR under different lighting conditions by testing CleAR on an augmented Laval dataset with diverse lighting conditions of light intensity and color temperatures. Qualitatively, our user study also confirms the effectiveness and robustness of CleAR, which shows at least 12% higher quality ratings.
@@
Related works on mobile AR lighting estimation systems seek to extract environment information from physical light probes (Prakash et al., 2019), user dynamics, and learning-based solutions (Zhao and Guo, 2021; Gardner et al., 2019; Wang et al., 2022; Yang et al., 2023). While physical light probes provide the most comprehensive environment observations, their use in practical applications is typically constrained due to the need for physical light probe presence. While environment lighting can also be extracted from dynamic environment observations, the estimation is usually incomplete as AR devices often do not have comprehensive environment observations. In contrast, leveraging learned models to estimate environment lighting from AR device camera images is a more feasible solution for AR applications. Over the past couple of decades, learning-based methods have evolved from discovering scene lighting cues from image details, such as highlights and shadows (Yu et al., 1999), to regress omnidirectional environment lighting representations (Zhao and Guo, 2020, 2021; Gardner et al., 2019). However, autoregressive models cannot effectively tackle the environment information generation in lighting estimation. For example, Xihe (Zhao and Guo, 2021) provides real-time low-frequency lighting estimation to AR applications but fails to support detailed environment reflections on object rendering. More recently, a new opportunity has arisen as generative models can support highly detailed image content for environment map estimation (Wang et al., 2022; Yang et al., 2023). In this work, we focus on novel AR system integrations with image-generative models to provide high-quality environment lighting estimation. We specifically focus on adapting generative lighting estimation models to achieve robust estimation under challenging environmental lighting conditions.
We summarize our main contribution as the following:
@@
  • We design and implement a novel AR context-guided generative lighting estimation system, CleAR. Our design leverages generative models to tackle the environmental observation limitation of AR devices. Specifically, our system uses a novel two-step generative lighting estimation pipeline to estimate an accurate environment map with visual details that match the AR user’s physical environment. To train the generative models, we craft a large-scale lighting estimation dataset with approximately 30K data items.
@@
  • We present a measurement study on recent lighting estimation datasets to understand the complexity of environmental lighting conditions and explore the challenges of achieving robust lighting estimation. With the insights from this study, we design and conduct lighting estimation robustness testing experiments to evaluate several lighting estimation systems, including ours.
@@
  • We develop two real-time estimation refinement techniques to improve the estimation quality of our system. Our first refinement component can automatically select the best environment map from generated results to match the current lighting conditions. Our second refinement component can match the color appearances between generated environment maps with real-time camera view to improve estimation robustness.
@@
  • We implement CleAR as an end-to-end edge-assisted framework that can be integrated into third-party AR applications. We demonstrate the integration with an example object placement-based AR application. We will provide the dataset, system source code, and demo application links once the paper is accepted for publication to encourage follow-up research.
  • We conduct comprehensive experiments to evaluate the effectiveness of CleAR. Specifically, we tested CleAR with SoTA lighting estimation systems on standard testing datasets and our robustness testing dataset. We also conduct a quality assessment study with human perception feedback on the lighting estimation results. Overall, the CleAR rating score ranked the top with 12% higher than the second best method. Additionally, the overall standard deviation of the rating values for CleAR is also lower than the second best method (1.67) by 7%.
@@ 2. Background
Lighting Estimation in Mobile AR. Environmental lighting provides important environmental information and plays a crucial role in creating photorealistic visual effects when compositing virtual objects into real-world backgrounds (Akimoto et al., 2022). Lighting estimation has been a long-standing research question in the vision and graphics community (Li et al., 2023). On mobile AR, estimating accurate environment lighting is important for creating immersive visual experiences when overlaying virtual content on physical environments. Such application includes virtual try-ons (Zhao et al., 2023) and interactive games (Scorpio et al., 2022). Obtaining accurate environment lighting on mobile AR faces several additional challenges compared to traditional image—or video-based lighting estimation tasks. For example, mobile AR devices usually have limited camera observation ranges, i.e., FoV. Because visually coherent AR experiences demand omnidirectional environment lighting, such camera observation capability limitation makes lighting estimation on mobile AR a highly ambiguous process and, thus, an ill-posed problem. Solving such problems typically requires a long computation time, while mobile AR applications demand timely environment lighting updates to maintain visual coherency in temporally changing lighting. To tackle these challenges, mobile system researchers typically aim to acquire environment information by leveraging spatial-temporal 3D environment observations (Zhao and Guo, 2021) or guided user movements (Zhao et al., 2022). While increased environmental observation positively impacts the accuracy of lighting estimation, it does not sufficiently ensure the robustness of the results.
@@
Conditional Generative Models. Recently, generative models have demonstrated impressive capabilities in high-quality data synthesis (Yang et al., 2023; Croitoru et al., 2023). Generative models have demonstrated impressive capabilities in high-quality data synthesis. Traditional generative models, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), Variational Autoencoders (VAEs) (Kingma, 2013), and autoregressive models (Van Den Oord et al., 2016), have demonstrated success in generating realistic images, text, and other types of data. However, these models are typically unconditional, meaning that they generate data without external guidance or specific control over certain attributes of the output. Conditional generative models extend generative models by incorporating additional information, referred to as conditions, to guide the generation process. These conditions can range from class labels in image generation (Mirza, 2014) to textual descriptions in text-to-image models (Reed et al., 2016) or even semantic features for structured data generation (Hong et al., 2018). In this work, we leverage ControlNet (Zhang et al., 2023), a state-of-the-art conditional diffusion model, to implement context-guided generative lighting estimation. ControlNet allows conditional image generation from several conditioning sources, such as text, incomplete images, semantic maps, depth maps, key points, and edge maps. Our research leverages ControlNet’s conditional capabilities to incorporate rich and diverse information embedded in the AR context. Specifically, we explore how partial environment RGB observation, environment semantic map, and ambient lighting conditions can be effectively integrated into ControlNet to enhance immersive AR experiences.
@@ 3. Motivation and Challenges
We study the challenges of robust mobile AR lighting estimation through a measurement study. In §3.1, we explore the complexity of environment lighting conditions and the pre-existing biases in existing lighting estimation datasets. Then, in §3.2, we formally summarize the challenges.
3.1. Lighting Condition Measurement
To understand the challenges in robust lighting estimation, we first conduct a measurement study of environmental lighting conditions on standard lighting estimation datasets. Our measurement focuses on two properties of environmental lighting conditions: light intensity and color temperature, representing an environment’s overall brightness and color appearances. We choose these two lighting condition properties because of their significant impact on virtual object rendering of all materials kinds (Zhao et al., 2022) We select two standard lighting estimation data sources: the the Laval dataset (Gardner et al., 2017), an academic open research dataset, and the PolyHaven website (pol, 2024), a royalty-free HDR environment map data website. In total, we collected 2235 and 196 data items from the Laval dataset and PolyHaven, respectively. Following (Phongthawee et al., 2023; Wang et al., 2022), we perform a standard data preprocessing procedure of a color correction process with gamma correction with $\gamma=2.4$ and setting the 99th percentile of pixel intensity to 0.9.
@@ 3.1. Lighting Condition Measurement
Figure 2. Lighting condition measurement. We measure the distributions of light intensity (2a) and color temperature (2b) on environment maps collected from the Laval indoor dataset and PolyHaven. While both datasets include diverse lighting conditions, we observe several distribution biases in the datasets. For example, a significant part of the PolyHaven data items lies in the cool color range. Also, both datasets have an imbalanced distribution between neutral and other lighting conditions.
where $i$ represents the i-th pixel in the environment map with $N$ pixels. For color temperature value extraction, we first calculate the average pixel RGB value from each environment map and then calculate its correlated color temperature using a recent method (Ohno, 2014). The calculated environment map image color temperature values are given in the Kelvin unit. In Figure 2, we visualize the measured lighting condition distributions and annotate the human-perceivable categories of the light intensity and color temperature conditions. While both datasets contain a wide selection of lighting conditions, the data distribution shows several biases. Predominantly, neutral lighting conditions of light intensity and color temperatures are seen in two datasets. For light intensity, more low-light environments than bright-lighting conditions data items are seen in the Laval dataset. As for the color temperature distribution, a significant part of the PolyHaven data items lies in the cool color range. These data distribution biases have also led to biased training and evaluation of lighting estimation systems. In particular, we will show robustness testing results of several recent lighting estimation models in §6.
3.2. Robust Lighting Estimation Challenges
Here, we summarize two key challenges in designing a robust mobile AR lighting estimation system.
3.2.1. Diverse Lighting Conditions.
As our measurement study suggests, real-world environment lighting has complex physical properties, such as light intensity and color temperatures. Achieving accurate results in diverse lighting conditions requires estimation systems to match the estimated environment map color profiles to the lighting conditions while maintaining the estimated environment image visual details. This requirement demands lighting estimation systems to have rich built-in knowledge of diverse environmental lighting conditions. Additionally, as mobile AR devices only have limited environment observation, estimating high-quality environment maps requires lighting estimation systems to complete an omnidirectional environment view based on existing observations. This inherently ambiguous process demands the system to create visual contexts that closely match the physical surroundings of the AR user, while accurately estimating high-intensity light sources to ensure realistic rendering of highlights and shadows on virtual objects. Moreover, real-world environmental lighting conditions often vary from scene to scene and from time to time. For example, indoor rooms with daytime sunlight usually appear in neutral color temperatures, while rooms lit with LED light sources can appear in either warm or cold colors. To ensure the consistency of mobile AR visual experiences, lighting estimation systems must also adapt their estimation results to changing environment lighting conditions in real-time.
3.2.2. Data Availability and Biases.
Our measurement study also shows that existing standard lighting estimation datasets do not have a balanced light intensity distribution or color temperatures. When used to train lighting estimation models, these data biases can lead to model generalization issues in underrepresented lighting conditions. Furthermore, the availability of high-quality HDR lighting estimation datasets is extremely limited. To capture a high-quality HDR environment map, researchers usually need to use a professional DSLR camera to take multiple images of the environment at different angles and under different camera exposures (Debevec, 2006). This process can be costly and time-consuming. Even in cases where professional setups can be created for lighting ground truth capturing, it is still impractical to capture a dataset that contains comprehensive lighting conditions and environmental details. Additionally, these data biases can also make it difficult to conduct fair evaluations to test the robustness of existing lighting estimation systems. As a result, current lighting estimations often overlook the importance of robustness under different lighting conditions and fail to adapt to the complex real-world environments of mobile AR users.

### v2

@@ Abstract.
High-quality environment lighting is essential for creating immersive mobile augmented reality (AR) experiences. However, achieving visually coherent estimation for mobile AR is challenging due to several key limitations in AR device sensing capabilities, including low camera FoV and limited pixel dynamic ranges. Recent advancements in generative AI, which can generate high-quality images from different types of prompts, including texts and images, present a potential solution for high-quality lighting estimation. Still, to effectively use generative image diffusion models, we must address two key limitations of content quality and slow inference. In this work, we design and implement a generative lighting estimation system called CleAR that can produce high-quality, diverse environment maps in the format of $360^{\circ}$ HDR images. Specifically, we design a two-step generation pipeline guided by AR environment context data to ensure the output aligns with the physical environment’s visual context and color appearance. To improve the estimation robustness under different lighting conditions, we design a real-time refinement component to adjust lighting estimation results on AR devices. To train and test our generative models, we curate a large-scale environment lighting estimation dataset with diverse lighting conditions. Through a combination of quantitative and qualitative evaluations, we show that CleAR outperforms state-of-the-art lighting estimation methods on both estimation accuracy, latency, and robustness, and is rated by 31 participants as producing better renderings for most virtual objects. For example, CleAR achieves 51% to 56% accuracy improvement on virtual object renderings across objects of three distinctive types of materials and reflective properties. CleAR produces lighting estimates of comparable or better quality in just 3.2 seconds—over 110X faster than state-of-the-art methods. Moreover, CleAR supports real-time refinement of lighting estimation results, ensuring robust and timely updates for AR applications.
@@ 1. Introduction
To address this challenging task, traditional systems often employ autoregressive models (Zhao and Guo, 2021) to extract parametric lighting information from AR device camera images (shown in Fig. 1a). While these methods effectively capture coarse environmental lighting conditions, they struggle to reliably capture high-frequency details, such as geometric features of the environment. To address the critical needs of high-frequency information in environment lighting, recent works (Somanath and Kurz, 2021; Yang et al., 2023; Wang et al., 2022) have sought new solutions by leveraging the advancements in controllable generative models. The primary reason is that these models have the potential to generate fine-grained environmental details (shown in Fig. 1a), thereby empowering AR systems with a more realistic rendering effect. However, it is challenging to achieve robust mobile AR lighting estimation with generative models in real-world environments. We identify two key challenges below.
@@
First, robust lighting estimation demands accurate and timely estimations. Therefore, it is crucial to ensure that the system can provide reliable and accurate estimation outputs under diverse and complex lighting conditions. Additionally, generative models often exhibit long inference latencies, which slow down the estimation process. Directly integrating generative model-based lighting estimation methods may not satisfy the needs of real-time mobile AR experiences. Second, training and evaluating robust lighting estimation models require careful attention to data bias. Specifically, the datasets used must be curated to reflect diverse and balanced lighting conditions to ensure fair and generalizable performance. However, our measurement study on existing datasets reveals distribution biases in key lighting properties, such as intensity and color temperature, which in turn affect the generalization and robustness of several recent lighting estimation models (Wang et al., 2022; Akimoto et al., 2022).
In this paper, we address the above issues with CleAR, a novel generative lighting estimation system for mobile AR. To support robust lighting estimation, CleAR builds on two key ideas: (i) AR context-guided generative estimation and (ii) edge-device collaborative estimation. It leverages multimodal AR context to guide 360 HDR environment map estimation using generative diffusion models. To ensure temporal consistency, CleAR uses a multi-output strategy to generate environment maps that best match real-time lighting conditions. To improve estimation quality and responsiveness, CleAR applies color appearance matching for fine-grained adjustments to environment maps.
@@
At the core of CleAR, the lighting estimation process in CleAR revolves around a two-step generative pipeline that estimates 360 HDR environment maps from partial LDR environment observations. Our key design insight is to separate the generative model training objective into two domains: LDR environment map completion and high-intensity pixel value estimation. This novel learning objective design addresses the practical challenge of the scarcity of high-quality lighting estimation training data, allowing us to leverage pre-trained large models to tackle each generation step effectively. Beyond its two-step generative pipeline, CleAR leverages AR context—semantic maps to guide visual details and ambient light sensor data to inform lighting intensity and color temperature. This context-guided generative estimation design effectively helps CleAR to tackle the challenge of estimating omnidirectional environment maps from limited environment observations on mobile AR devices.
@@
To evaluate CleAR, we conduct a comprehensive evaluation that includes in-lab deployment tests, data-driven evaluations, and a user study. In the deployment test, CleAR excels in supporting virtual object rendering quality compared to three representative baselines: unwrapping a mirror ball (physical reference) (Debevec, 2006), ARKit (commercial) (Apple, 2022), and LitAR (academic) (Zhao et al., 2022a). Quantitatively, we compare CleAR with state-of-the-art lighting estimation models (Phongthawee et al., 2023; Wang et al., 2022; Akimoto et al., 2022; Gardner et al., 2017) and show that CleAR outperforms the best performing baseline, DiffusionLight (Phongthawee et al., 2023), by up to 56%. Our user study also confirms the effectiveness and robustness of CleAR, showing at least a 12% improvement in quality ratings compared to StyleLight, the second-best method. Furthermore, we introduce a robustness testing protocol, which leverages our augmented Laval dataset to test the estimation accuracy under diverse lighting intensity and temperature conditions. We observe that CleAR consistently achieves low estimation errors across diverse lighting conditions, whereas the estimation quality of baseline methods varies significantly with changes in lighting.
@@
Related works on mobile AR lighting estimation systems seek to extract environment information from physical light probes (Prakash et al., 2019), user dynamics, and learning-based solutions (Zhao and Guo, 2021; Gardner et al., 2019; Wang et al., 2022; Yang et al., 2023). While physical light probes provide the most comprehensive environment observations, their use is limited by the need for physical setup. Consequently, AR applications often use learned models to estimate lighting from camera images. Over the past couple of decades, learning-based methods have evolved from discovering scene lighting cues from image details, such as highlights and shadows (Yu et al., 1999), to regressing omnidirectional environment lighting representations (Zhao and Guo, 2020, 2021; Gardner et al., 2019). However, autoregressive models cannot effectively tackle the environment information generation in lighting estimation. For example, Xihe (Zhao and Guo, 2021) provides real-time low-frequency lighting estimation for AR applications but lacks support for detailed environment reflections in object rendering. In contrast, recent advances in generative models enable the estimation of highly detailed environment maps (Wang et al., 2022; Yang et al., 2023), opening new possibilities for visually coherent rendering. Our work explores the integration of image-generative models into AR systems to enable high-quality environment lighting estimation, with novel AR context-aware design to improve generation speed, accuracy, and robustness. We summarize our main contribution as follows:
@@
  • We design and implement CleAR, a novel AR context-guided generative lighting estimation system that overcomes the limited environmental sensing of AR devices. CleAR uses a two-step generative pipeline to produce accurate, visually detailed environment maps aligned with the user’s physical surroundings. We also create a large-scale lighting estimation dataset with $\sim$30K data items for training generative models.
@@
  • We conduct a measurement study of recent lighting estimation datasets to analyze environmental lighting complexity and identify challenges to robustness. Guided by these insights, we design robustness testing experiments to evaluate several lighting estimation systems, including CleAR.
@@
  • We develop a real-time refinement technique to enhance estimation quality. The refinement works by first matches the color appearance between the generated environment map and the real-time camera view to improve robustness, and then select the best environment map to match current lighting.
@@
  • We implement CleAR as an end-to-end, edge-assisted framework that can be integrated into third-party AR applications. To demonstrate its usability, we integrate it into an example AR app for virtual object placement.
  • We evaluate CleAR through both quantitative and qualitative experiments, comparing it against SoTA lighting estimation systems and models on a commonly used dataset (Gardner et al., 2017) and a robustness-focused variant generated by us. A user study based on perceptual quality ratings shows that CleAR outperforms all baselines, scoring 12% higher than the second-best method, with 7% lower rating variance.
@@ 2. Background
Environment Lighting Representations. Lighting estimation has been a long-standing research question in the vision and graphics community (Li et al., 2023). Pioneering works have established several methods for capturing high-fidelity lighting of physical environments. For instance, environmental lighting can be captured with mirror balls, panoramic cameras, and photometric stereo techniques (Debevec, 2006; Ward et al., 2008; Vlasic et al., 2009). The captured environment lighting can also be represented in various formats, including parametric format (Green, 2003), neural representations (Pandey et al., 2021), and image-based format (Debevec, 2006). In modern computer graphics, mirror ball-based capturing and image-based lighting are widely used for accurate lighting rendering. However, mirror ball setups can be cumbersome in practice (Prakash et al., 2019), making image-based lighting a more practical approach for mobile AR, which this work adopts. Image-based lighting relies on acquiring high-dynamic range (HDR) environment maps that capture omnidirectional lighting at a specific location. HDR images are critical because they preserve the full range of luminance—including bright highlights and subtle shadows—often lost in standard low-dynamic range (LDR) images (Dufaux et al., 2016). To enable realistic lighting effects and photorealistic virtual content, this work focuses on generating HDR environment maps.
@@
Lighting Estimation in Mobile AR. In mobile AR, accurately estimating environment lighting is crucial for creating immersive visual experiences when overlaying virtual content on physical environments. Such application includes virtual try-ons (Zhao et al., 2023) and interactive games (Scorpio et al., 2022). Obtaining accurate environmental lighting in mobile AR faces several additional challenges compared to traditional image— or video-based lighting estimation tasks. For example, mobile AR devices typically have limited environmental sensing capabilities, such as the field of view (FoV) of their cameras. Because visually coherent AR experiences require omnidirectional environment lighting, the limitations in environment sensing render lighting estimation in mobile AR a highly ambiguous process and, thus, an ill-posed problem. Traditional methods (Song and Funkhouser, 2019; Gardner et al., 2017, 2019) often aim to solve this problem using deep models, which result in lengthy computation times. However, mobile AR applications require timely environment lighting updates to maintain visual coherence in temporally changing lighting environments. Recently, mobile system researchers have begun leveraging enhanced device capabilities (Zhao and Guo, 2021) and user-in-the-loop interactions (Zhao et al., 2022a) to acquire more reliable environmental information, thereby improving the accuracy and robustness of lighting estimation in mobile AR. Despite these advances, current state-of-the-art methods still fall short of providing the accuracy and robustness required for reliable lighting estimation in mobile AR applications.
@@
Conditional Generative Models. Recent years have witnessed significant advancements in the data synthesis capabilities of generative models (Yang et al., 2023; Croitoru et al., 2023). Various model architectures, such as Generative Adversarial Networks (GANs) (Goodfellow et al., 2020), Variational Autoencoders (VAEs) (Kingma, 2013), have achieved notable success in generating realistic images, text, and other forms of data. However, the content generation process in these models typically cannot be conditioned. In other words, users of these models cannot control the outputs through external guidance or specific input signals. Conditional generative models extend generative models by incorporating additional input information, referred to as conditions, to guide the generation process. These conditions can range from class labels in image generation (Mirza, 2014) to textual descriptions in text-to-image models (Reed et al., 2016) or even semantic features for structured data generation (Hong et al., 2018). In this work, we base CleAR on the ControlNet (Zhang et al., 2023) architecture, a type of state-of-the-art conditional diffusion model. ControlNet works by extending a pre-trained diffusion model, such as Stable Diffusion (Rombach et al., 2021), with additional neural network layers that allow external conditions to influence the diffusion process. Specifically, ControlNet uses a trainable copy of the diffusion model’s encoder to encode the input conditioning latent code and merge it with the diffusion model’s latent code. By doing so, it maintains the strong generative priors of the base diffusion model while allowing the generation process to be guided by external conditioning inputs.
@@ 3. Motivation and Challenges
This section presents a measurement study of two key lighting attributes—light intensity and color temperature—to identify challenges for robust lighting estimation systems. Both attributes are essential for photorealistic and consistent rendering across diverse material types in mobile AR (Zhao et al., 2022a).
Our analysis uses two standard lighting estimation datasets: the Laval dataset (Gardner et al., 2017), an academic open research dataset, and the PolyHaven website (pol, 2024), a royalty-free HDR environment map data website. In total, we collected 2235 and 196 data items from the Laval dataset and PolyHaven, respectively. On these data, we calculate the light intensity as the total luminance of a given HDR environment map image, along with the environment map image’s color temperature values in Kelvin. Additional details of analysis setup, measurement calculations, and technical details are included in Appendix §A. Figure 2 shows the lighting condition distributions with annotated human-perceivable categories for light intensity and color temperature.
We make the following key observations. First, both datasets include a wide range of light intensity and color temperature values. In particular, the measured light intensity spans from as low as 0.6 to over 8 in luminance values, capturing both dim and brightly lit scenes. Similarly, the color temperature varies widely, ranging from below 1000K, typical of incandescent lighting, to over 9000K, which is seen in artificial lights. This suggests that real-world lighting conditions have complex and diverse properties. Second, there is a notable imbalance in the distribution of lighting conditions across the datasets. Specifically, both datasets are skewed toward neutral light intensity and color temperature values. The Laval dataset contains more low-light environment samples than bright-light ones. Meanwhile, a substantial portion of the PolyHaven dataset’s color temperature values fall within the cooler range.
@@
Figure 2. Distributions of lighting condition in two datasets. We analyze the distributions of light intensity and color temperature on environment maps collected from the Laval indoor dataset and PolyHaven. Both datasets exhibit data biases. For example, a significant part of the PolyHaven data items lies in the cool color range. Also, both datasets have an imbalanced distribution between neutral and other lighting conditions.
Our analysis reveals a critical challenge for achieving robust lighting estimation: complex and diverse real-world lighting conditions. In many real-world scenarios, environmental lighting conditions often vary across scenes and over time. For example, indoor rooms lit by sunlight often exhibit neutral to warm color temperatures, while those illuminated by LED sources tend to appear cooler. Therefore, to achieve robust estimation accuracy, sophisticated system designs are needed to understand and adapt to diverse environmental lighting conditions. In particular, lighting estimation systems must have rich built-in knowledge of diverse environmental lighting conditions, as well as adaptive application-time policies to react to complex AR device sensor outputs and user movements. In other words, lighting systems should be capable of learning generalizable lighting representations and not be overly sensitive to changes in environmental lighting. To address this challenge, in §4, we introduce an end-to-end system that takes multimodal AR data to guide the generation of a visually coherent HDR environment map with color refinement and color palette matching.
Another key challenge revealed by our study is the inherent data biases, which pose challenges to training and evaluating generative models in lighting estimation systems. These biases in training data can lead to overfitting on dominant lighting conditions and poor generalization to underrepresented lighting features. For instance, models trained predominantly on neutral or cool lighting conditions may fail to produce plausible estimations under warm or high-intensity illumination. Furthermore, biased data distributions in testing data can skew evaluation results, making it difficult to evaluate the robustness of lighting estimation systems. To address the data bias issue, We introduce a data balancing technique in Appendix §B, and present a robustness testing protocol in §6, through which we evaluate several recent lighting estimation models.